OpenStudy (anonymous):

What is the solution to the following equation? 2(3x − 7) + 18 = 10 1 3 5 6

OpenStudy (anonymous):

@DIDITHEDAD @joojoe

OpenStudy (anonymous):

distribute again

OpenStudy (anonymous):

2 * 3x 2 * -7

OpenStudy (anonymous):

6x

OpenStudy (anonymous):

-14

OpenStudy (anonymous):

Yuup! 6x - 14 + 18 = 10

OpenStudy (anonymous):

Add together 14 and 18

OpenStudy (anonymous):

lol 32

OpenStudy (anonymous):

6x - 32 = 10 now add 32 (because it is being subtracted on the left side, we do teh opposite which is add.) so...6x - 32 + 32 = 10 + 32 6x = 10 + 32 what is 10 + 32

OpenStudy (anonymous):

42

OpenStudy (anonymous):

6x = 42 divide both sides 42/6 is what?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

Hang on we did something wrong...

OpenStudy (anonymous):

thats my fav. number btw

OpenStudy (anonymous):

I see what we did....instead of what is 14 + 18 what is 18 - 14

OpenStudy (anonymous):

4

OpenStudy (anonymous):

subtract 10 - 4

OpenStudy (anonymous):

6

OpenStudy (anonymous):

6x = 6 divide by 6 x = 1 answer is 1

OpenStudy (anonymous):

You getting the hang of it?

OpenStudy (anonymous):

yep lol sorry for the delay i had thouhgt i sent it thx so much i have another one

OpenStudy (anonymous):

GREAT!